Shift report 20030122 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Jan 22 2003 - 04:00:47 EST

  • Next message: Ian Bearden: "TDC calibration"
    ------------------------------------------------------------
    Date: Wed Jan 22 04:00:47 2003, Shift: 00:00-08:00
    Supervisor: JIJ, with: IA
    Log book#13 pages 118 - 119
    
    Shift Summary:
    ------------------------------------------------------------
    1.
    No beam. Some attempted ramps with 110 bunches, but beam was lost around transition every time. Then 56 bunches, but it did not seem to change much. Some magnet problem which was solved quickly.
    
    2.
    7542 - Pedestal run
    7543 - TPC pedestal run
    7544 - DVM run
    
    3.
    None detected.
    
    Machine study from 4 a.m. Therefore we sign off early and go to bed.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Wed Jan 22 2003 - 04:01:28 EST